myininaya (myininaya):

calculus dream problem? (fg)'=f'g' and (fg)''=f''g'' I don't think this what you guys are looking for but here what i did

OpenStudy (watchmath):

again? :D

OpenStudy (anonymous):

watchmath: ur problem has become famous..:D

myininaya (myininaya):

so this is your problem?

myininaya (myininaya):

have you done it?

OpenStudy (watchmath):

Check it here. It is solution for f. The g will be of the same form, but then the coefficients need to be adjusted.

myininaya (myininaya):

i had a dream about this problem last night

OpenStudy (watchmath):

:D :D

myininaya (myininaya):

i feel like i really dont sleep when i dream about math problems

myininaya (myininaya):

i have to look at this later but im very interested thanks for the link

myininaya (myininaya):

pdf*

OpenStudy (watchmath):

you are welcome :)

OpenStudy (anonymous):

i got as far as g = C e^ int ( f' / ( f' - f)

OpenStudy (anonymous):

i think, why did someone say , let F = e ^f, and G = e^g

OpenStudy (anonymous):

then ln F = f , and ln G = f ,

OpenStudy (watchmath):

Yes, that's good laplace. Then basically the idea is that from that g we can express, g' and g'' in terms of g. By doing that, then we have an equation in terms of f,f' and f'' only.

OpenStudy (watchmath):

I mean after you find \(g=Ce^{\int f'/(f'-f)}\)

OpenStudy (watchmath):

Laplace, are you cantorset?

OpenStudy (anonymous):

yes, but dont tell anyone

OpenStudy (watchmath):

:D :D

OpenStudy (anonymous):

my account has been suspended,

OpenStudy (anonymous):

im here to do math, not to debate with admin :)

OpenStudy (anonymous):

so from that g , whats g ' , its ...

OpenStudy (watchmath):

do you see the pdf file that I attached above?

OpenStudy (anonymous):

g ' = f' ( f' - f) C e ^ ( int f' / f' - f ) ?

OpenStudy (watchmath):

correct, which is f'g/(f'-f)

OpenStudy (anonymous):

oh im going backwards,

OpenStudy (anonymous):

this problem has many twists and turns

OpenStudy (anonymous):

one sec , im going to read

OpenStudy (watchmath):

there is some little mistake whenn I said that if f'=f then f'g=0 it should be then fg'=0. But similar to what I wrote we will arrive a t contradiction.

OpenStudy (anonymous):

you can fix it and send me new pdf?

OpenStudy (anonymous):

or you already updated? i dont know how to write pdf's, thats pretty cool

OpenStudy (watchmath):

hold on

OpenStudy (anonymous):

i also want to solve that torus problem,

OpenStudy (watchmath):

Here is the updated version:

OpenStudy (watchmath):

Are you teaching as a TA laplace?

OpenStudy (anonymous):

no, i tutor at the moment online

OpenStudy (anonymous):

a TA is what? a teacher assistant?

OpenStudy (watchmath):

Yes, usually if you take a graduate program in math you are a TA as well.

OpenStudy (anonymous):

i should,

OpenStudy (anonymous):

sorry was away, i enjoy math. but sometimes i cant remember the last problem i was working on. some people say its ok to forget all the loose ends problems, and start fresh. dunno. i have various interests in math, linear algebra, diff eq. they always end up in an interesting problem

OpenStudy (watchmath):

If you have an interesting problem please post it here: http://www.ask.watchmath.com

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

whats an easy way to draw a graph and upload it? microsoft paint?

OpenStudy (watchmath):

is it a graph of a well defined function? if it is you can plot it at http://fooplot.com and save it as an image

OpenStudy (anonymous):

and how would you upload it, whats a good site

OpenStudy (anonymous):

i use sendspace.com, not sure how long they keep it up

OpenStudy (watchmath):

where do you want to upload the picture?

OpenStudy (anonymous):

well i can draw problems and solve them, and you can look over

OpenStudy (anonymous):

www.sendspace.com , is a website that you can upload anything

OpenStudy (anonymous):

ok heres a question i was working on, i got no solution 2/ln 3 * ln ( x - 5) = log base 3 ( x-2) + log base 3 (5x) , i made the right side a product and then changed base

OpenStudy (watchmath):

so the left hand side is \(\frac{2}{\ln 3\ln(x-5)}\)

OpenStudy (anonymous):

nope, ln (x-5) in numerator

OpenStudy (anonymous):

i know, its ambiguous when you have a bad student :)

OpenStudy (anonymous):

im careful, i use CAS (computer algebra software)

OpenStudy (anonymous):

maple, did you know that maple cant find tan (105 degrees) ,

OpenStudy (anonymous):

tan ( 105* Pi/180) , its an exact angle, i want to complain to the makers

OpenStudy (watchmath):

Oh ok, then by the change of base formula we have \(2\log_3(x-5)=\log_3((x-2)(5x))\) So \((x-5)^2=(x-2)(5x)\) \(x^2-10x+255=5x^2-10x\) \(4x^2=225\) \(x=\pm\frac{15}{2}\) only the postive one works

OpenStudy (watchmath):

sorry ..... It should be 4x^2=25 So x=5/2 then no solution as well

OpenStudy (anonymous):

hah, yeah

OpenStudy (anonymous):

weird to give a student

OpenStudy (anonymous):

you changed the left side to log base 3

OpenStudy (anonymous):

2/ln 3 * ln ( x - 5) = log base 3 ( x-2) + log base 3 (5x) , what happened to your left side?

OpenStudy (anonymous):

your left side is log _3 (x-5) / log _3 ln 3 * 2/ln 3

OpenStudy (watchmath):

remember that \[\log_ba=\frac{\log_c a}{\log_cb}\] Where you can choose your c as you like. Choosince c to be the natural number e, then the left side is exactly \(\log_3((x-5))\)

OpenStudy (anonymous):

right, i followed the rule

OpenStudy (anonymous):

but you have 2/ln 3 *log _3 (x-5) / log _3 (ln 3)

OpenStudy (anonymous):

and whats log _3 of ln 3 ? log base 3 of ln 3

OpenStudy (watchmath):

In that case I thin we can't simplify it nicely.

OpenStudy (anonymous):

its funny you did that and still got right answer. my instinct would be to change the right side to ln instead of log base 3

OpenStudy (anonymous):

so it must be the case that ln 3 * log_3 ln 3 = 1 ?

OpenStudy (watchmath):

no

OpenStudy (watchmath):

it is equal to ln(ln 3)

OpenStudy (anonymous):

oh i think i see it now

OpenStudy (anonymous):

for the left side we have 2/( ln 3) * ln (x-5) = 2/ln 3 * log_3 (x-5) / log _3 (e) = 2 *log_3 (x-5) * 1/ ( ln 3 * log _3 (e) = log _3 (x-5)^2 * 1/ ( log _3 e^(ln 3) ) = log _3 (x-5)^2* 1 / ( log _3 3 ) = log_3 (x-5)^2 * 1